OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (anonymous):

Use the point slope formula: \[m=4\] \[x _{0} = -3\] \[y _{0}= -1\] \[y-y _{0}=m(x-x _{0})\]

OpenStudy (anonymous):

So: y - (-1) = 4 (x - (-3)) y + 1 = 4x + 12 y = 4x + 11